# Partial information decomposition as information bottleneck

### v1

@@ Abstract
The partial information decomposition (PID) aims to quantify the amount of redundant information that a set of sources provide about a target. Here we show that this goal can be formulated as a type of information bottleneck (IB) problem, which we term the “redundancy bottleneck” (RB). The RB formalizes a tradeoff between prediction and compression: it extracts information from the sources that predicts the target, without revealing which source provided the information. It can be understood as a generalization “Blackwell redundancy”, which we previously proposed as a principled measure of PID redundancy. The “RB curve” quantifies the prediction/compression tradeoff at multiple scales. This curve can also be quantified for individual sources, allowing subsets of redundant sources to be identified without combinatorial optimization. We provide an efficient iterative algorithm for computing the RB curve.
@@ I Introduction
Many research fields that study complex systems are faced with multivariate probabilistic models and high-dimensional datasets. Prototypical examples include brain imaging data in neuroscience, gene expression data in biology, and neural networks in machine learning. In response, various information-theoretic frameworks have been developed in order to study such multivariate systems in a universal manner. Here we focus on two such frameworks, the partial information decomposition and the information bottleneck.
@@
The partial information decomposition (PID) considers how information about a target random variable $Y$ is distributed among a set of source random variables $X_{1},\dots,X_{n}$ [1, 2, 3, 4]. For example, in neuroscience, the sources $X_{1},\dots,X_{n}$ might represent the activity of $n$ different brain regions and $Y$ might represent a stimulus, and one may wish to understand how information about the stimulus is encoded in different brain regions. A central idea of the PID is that the information provided by the sources can exhibit redundancy, when the same information about $Y$ is present in each source, and synergy, when information about $Y$ is found only in the collective outcome of all sources. Moreover, it has been shown that standard information-theoretic quantities, such as entropy and mutual information, are not sufficient to quantify redundancy and synergy, necessitating the development of new measures [1, 5]. However, finding the right measures of redundancy and synergy has proved difficult. In recent work [4], we showed that these measures can be naturally defined by formalizing the analogy between set theory and information theory that lies at the heart of the PID [5]. We then proposed a measure of redundant information (Blackwell redundancy) that is motivated by algebraic, axiomatic, and operational considerations. We argued that Blackwell redundancy overcomes many limitations of previous proposals [4].
@@
The information bottleneck (IB) [6, 7] is a method for extracting compressed information from one random variable $X$ that optimally predicts another target random variable $Y$. For instance, in the neuroscience example with stimulus $Y$ and brain activity $X$, the IB method could be used to quantify how well the stimulus can be predicted using only one bit of information about brain activity. The overall tradeoff between prediction of $Y$ and compression of $X$ is captured by the so-called IB curve. The IB method has been employed in various domains, including neuroscience [8], biology [9], and cognitive science [10]. In recent times, it has become particularly popular in machine learning applications [11, 12, 13, 14, 7].
@@
In this paper, we demonstrate a formal connection between PID and IB. We focus in particular on the relationship between IB and PID redundancy, leaving the connection to other PID measures (such as synergy) for future work. To begin, we show that Blackwell redundancy can be formulated as an information-theoretic optimization problem. This optimization problem extracts information from the sources that best predict the target, without revealing which source provided the information. We then define a generalized measure of Blackwell redundancy by relaxing the constraint; that is, we ask how much predictive information can be extracted from the sources without revealing more than a certain number of bits about the identity of the source. Our generalization leads to an IB-type tradeoff between prediction of the target (generalized redundancy) and compression cost (leakage of information about the identity of the source). We refer to this formulation as the redundancy bottleneck (RB), and to the resulting tradeoff between prediction and compression as the RB curve. We also show that the RB prediction and compression terms can decomposed into contributions from individual sources, giving rise to an individual RB curve for each source.
@@
Besides the intrinsic theoretical interest of unifying PID and IB, our approach brings about several practical advantages. In particular, the RB curve offers a fine-grained analysis on PID redundancy, showing how redundant information appears at various scales and across different sources. This fine-grained analysis can be used to uncover sets of redundant sources without performing intractable combinatorial optimization. Our approach has numerical advantages. The original formulation of Blackwell redundancy was based on a difficult optimization problem that becomes infeasible for larger systems. By reformulating Blackwell redundancy as an IB-type problem, we are able to solve it efficiently using an iterative algorithm, even for larger systems (code available at https://github.com/artemyk/pid-as-ib). Finally, the RB brings about has some attractive formal properties. For instance, unlike the original Blackwell redundancy, the RB curve is continuous in the underlying probability distributions.
@@
This paper is organized as follows. In the next section, we provide background on IB, PID, and Blackwell redundancy. In Section III, we introduce the RB, illustrate it with several examples, and discuss its formal properties. In Section IV, we introduce an iterative algorithm to solve the RB problem. We discuss implications and possible future directions in Section V. All proofs are found in the appendix.
@@ II Background
We begin by providing relevant background on information bottleneck, partial information decomposition, and Blackwell redundancy.
@@ II.1 Information bottleneck (IB)
The information bottleneck (IB) method provides a way to extract information that is present in one random variable $X$ that is relevant for predicting another target random variable $Y$ [6, 15, 16]. To do so, IB posits a “bottleneck variable” $Q$ that obeys the Markov condition $Q-X-Y$. This Markov condition guarantees that $Q$ does not contain any information about $Y$ that is not found in $X$. The quality of any particular choice of bottleneck variable $Q$ is quantified via two mutual information terms: $I(X;Q)$, which measures how well it compresses $X$, and $I(Y;Q)$, which measures how well it predicts $Y$. The IB method chooses $Q$ to maximize prediction given a constraint on compression [16, 15, 17],
$I_{\text{IB}}(R)=\sup_{Q:Q-X-Y}I(Y;Q)\quad\textrm{where}\quad I(X;Q)\leq R.$ (1)
@@
In practice, the IB curve is usually explored by considering the Lagrangian relaxation of the constrained optimization problem (1),
$F_{\text{IB}}(\beta):=\sup_{Q}I(Y;Q)-\frac{1}{\beta}I(X;Q)$ (2)
Here $\beta\geq 0$ is a parameter that controls the trade-off between compression cost (favored for $\beta\to 0$) and prediction benefit (favored for $\beta\to\infty$). The advantage of the Lagrangian formulation is that it avoids the non-linear constraint in Eq. (1). If the IB curve is strictly concave, then the two formulations (1) and (2) are equivalent, meaning that there is a one-to-one map between the solutions of both problems. When the IB curve is not strictly concave, a modified objective such as the “squared Lagrangian” or “exponential Lagrangian” should be used instead; see Refs. [18, 19, 20] for more details.
Since the original proposal, many reformulations, generalizations, and variants of the IB have been developed [7]. Notable examples include the “conditional entropy bottleneck” (CEB) [13, 21], the “multi-view IB” [22], the “distributed IB” [23], as well as a large family of objectives called the “multivariate IB” [24]. In general, all of these approaches consider a tradeoff between two information-theoretic terms — one term that quantifies prediction of target information that should be maximized, and one term that quantifies compression of unwanted information that should be minimized. We refer to an optimization that involves a tradeoff between information-theoretic prediction and compression terms as an IB-type problem.
@@ II.2 Partial information decomposition
The PID considers how information about a target random variable $Y$ is distributed across a set of source random variables $X_{1},\dots,X_{n}$. One of the main goals of the PID is to quantify redundancy, the amount of shared information that is found in each of the sources. The notion of redundancy in PID was inspired by an analogy between set theory and information theory that has re-appeared in various forms throughout the history of information theory [25, 26, 27, 28, 29, 30, 31]. If the amount of information provided by each source $X_{s}$ is conceptualized as the size of a set, then the redundancy is conceptualized as the size of the intersection of those sets [1, 5, 4]. Until recently, however, this analogy was treated mostly as an informal source of intuition, not as a rigorous technique.
In a recent paper [4], we demonstrated that the terms of the PID can be defined by formalizing this analogy to set theory. Recall that in set theory, the intersection of sets $A_{1},\dots,A_{n}$ is defined as the largest set $B$ that is contained in each set $A_{s}$ for $s\in\{1...n\}$. Thus, the size of the intersection of sets $A_{1},\dots,A_{n}$ is
$\left|\bigcap_{s=1}^{n}A_{s}\right|=\sup_{B}|B|\quad\textrm{where}\quad B\subseteq A_{s}\quad\forall s\in\{1\dots n\}.$
We showed that PID redundancy can be defined in a similar way. Formally, the redundancy between sources $X_{1},\dots,X_{n}$ is the maximum mutual information in any random variable $Q$ that is less informative about the target $Y$ than each individual source [4],
$I_{\cap}^{\sqsubseteq}:=\sup_{Q}I(Q;Y)\quad\textrm{where}\quad Q\sqsubseteq X_{s}\quad\forall s\in\{1\dots n\}.$ (3)
@@
We used a similar approach to define “union information”, which in turn leads to a principled measure of synergy [4]. Note that union information and redundancy are related algebraically but not numerically; in particular, unlike in set theory, the principle of inclusion-exclusion does not always hold. As mentioned above, here we focus entirely on redundancy, and leave the exploration of connections between IB and union information/synergy for future work.
@@ II.3 Blackwell redundancy
The Blackwell order is a relation over channels, i.e., conditional distributions with full support. Formally, some channel $\kappa_{B|Y}$ is said to be less informative than $\kappa_{C|Y}$ in the sense of the Blackwell order if there exists some other channel $\kappa_{B|C}$ such that
@@
Throughout, we use the notation $\circ$ to indicate composition of channels, as defined via matrix multiplication (for instance, $\kappa_{B|Y}=\kappa_{B|C}\circ\kappa_{C|Y}$ is equivalent to the statement $\kappa_{B|Y}(b|y)=\sum_{c}\kappa_{B|C}(b|c)\kappa_{C|Y}(c|y)$ for all $b$ and $y$). Eq. (4) In other words, $\kappa_{B|Y}$ is less informative than $\kappa_{C|Y}$ when $\kappa_{B|Y}$ can be produced by downstream stochastic processing of the output of channel $\kappa_{C|Y}$. We use the notation
@@
to indicate that $\kappa_{B|Y}$ is less Blackwell informative than $\kappa_{C|Y}$. The Blackwell order can also be defined over random variables rather than channels. Given a target random variable $Y$ with full support, random variable $B$ is said to be less Blackwell informative than random variable $C$ about $Y$, written as
@@
when their corresponding conditional distributions obey the Blackwell relation, $p_{B|Y}\preceq p_{C|Y}$ [36].
@@
The Blackwell order plays a fundamental role in statistics, and it has a important operational characterization in decision theory [37, 36, 38]. Specifically, $p_{B|Y}\preceq p_{C|Y}$ if and only if having access to channel $p_{C|Y}$ is better for every decision problem than having access to channel $p_{B|Y}$. See Refs. [39, 4] for details of this operational characterization and Refs. [36, 39, 40, 4, 41, 42] for more discussion of the relation between the Blackwell order and the PID.
@@
Combining the Blackwell order (6) with Eq. (3) gives rise to Blackwell redundancy [4]. Blackwell redundancy, indicated here as $I_{\cap}$, is the maximal mutual information in any random variable that is less Blackwell informative than each of the sources,
$I_{\cap}:=\sup_{Q}\,I(Q;Y)\quad\textrm{where}\quad Q\preceq_{Y}X_{s}\;\forall s.$ (7)
@@
$I_{\cap}$ has many attractive features as a measure of PID redundancy, and it overcomes many problems with previous approaches [4]. For instance, it can be defined for any number of sources, it satisfies the PID axioms, and it has fundamental statistical and operational interpretation. Statistically, it is the maximum information transmitted across any channel that can be produced by downstream processing of any one of the sources. Operationally, it is the maximum information that one can have about $Y$ without being able to do better on any decision problem than another agent that has access to one of the sources.
@@
The optimization problem (7) can be formulated as the maximization of a convex function subject to a set of linear constraints [4]. For a finite-dimensional system, the feasible set is a finite-dimensional polytope, and the maximum will lie on one of its extreme points, therefore the optimization can be solved exactly by enumerating the vertices of the feasible set and choosing the best one [4]. However, this approach is limited to small systems, because the number of vertices of the feasible set can grow exponentially.
Finally, observe that the target random variable $Y$ and the source random variables $X_{1},\dots,X_{n}$ enter into the definition of Blackwell redundancy only via the marginal distribution $p_{Y}$ and the channels $p_{X_{1}|Y},\dots,p_{X_{n}|Y}$. In particular, the joint distribution over $X_{1},\dots,X_{n}$ is not used and, in principle, may not even be defined. For this reason, it is more appropriate to consider Blackwell redundancy as a measure of redundancy in the channels $p_{X_{1}|Y},\dots,p_{X_{n}|Y}$, rather than as a measure of redundancy in the random variables $X_{1},\dots,X_{n}$. In the following, we generally work with the channel-based perspective.

### v2

# Partial information decomposition: redundancy as information bottleneck

@@ Abstract
The partial information decomposition (PID) aims to quantify the amount of redundant information that a set of sources provides about a target. Here, we show that this goal can be formulated as a type of information bottleneck (IB) problem, termed the “redundancy bottleneck” (RB). The RB formalizes a tradeoff between prediction and compression: it extracts information from the sources that best predict the target, without revealing which source provided the information. It can be understood as a generalization of “Blackwell redundancy”, which we previously proposed as a principled measure of PID redundancy. The “RB curve” quantifies the prediction–compression tradeoff at multiple scales. This curve can also be quantified for individual sources, allowing subsets of redundant sources to be identified without combinatorial optimization. We provide an efficient iterative algorithm for computing the RB curve.
@@ I Introduction
Many research fields that study complex systems are faced with multivariate probabilistic models and high-dimensional datasets. Prototypical examples include brain imaging data in neuroscience, gene expression data in biology, and neural networks in machine learning. In response, various information-theoretic frameworks have been developed in order to study multivariate systems in a universal manner. Here, we focus on two such frameworks, partial information decomposition and the information bottleneck.
@@
The partial information decomposition (PID) considers how information about a target random variable $Y$ is distributed among a set of source random variables $X_{1},\dots,X_{n}$ [1, 2, 3, 4]. For example, in neuroscience, the sources $X_{1},\dots,X_{n}$ might represent the activity of $n$ different brain regions and $Y$ might represent a stimulus, and one may wish to understand how information about the stimulus is encoded in different brain regions. A central idea of the PID is that the information provided by the sources can exhibit redundancy, when the same information about $Y$ is present in each source, and synergy, when information about $Y$ is found only in the collective outcome of all sources. Moreover, it has been shown that standard information-theoretic quantities, such as entropy and mutual information, are not sufficient to quantify redundancy and synergy [1, 5]. However, finding the right measures of redundancy and synergy has proven difficult. In recent work [4], we showed that such measures can be naturally defined by formalizing the analogy between set theory and information theory that lies at the heart of the PID [5]. We then proposed a measure of redundant information (Blackwell redundancy) that is motivated by algebraic, axiomatic, and operational considerations. We argued that Blackwell redundancy overcomes many limitations of previous proposals [4].
@@
The information bottleneck (IB) [6, 7] is a method for extracting compressed information from one random variable $X$ that optimally predicts another target random variable $Y$. For instance, in the neuroscience example with stimulus $Y$ and brain activity $X$, the IB method could be used to quantify how well the stimulus can be predicted using only one bit of information about brain activity. The overall tradeoff between the prediction of $Y$ and compression of $X$ is captured by the so-called IB curve. The IB method has been employed in various domains, including neuroscience [8], biology [9], and cognitive science [10]. In recent times, it has become particularly popular in machine learning applications [11, 12, 13, 14, 7].
@@
In this paper, we demonstrate a formal connection between PID and IB. We focus in particular on the relationship between the IB and PID redundancy, leaving the connection to other PID measures (such as synergy) for future work. To begin, we show that Blackwell redundancy can be formulated as an information-theoretic constrained optimization problem. This optimization problem extracts information from the sources that best predict the target, under the constraint that the solution does not reveal which source provided the information. We then define a generalized measure of Blackwell redundancy by relaxing the constraint. Specifically, we ask how much predictive information can be extracted from the sources without revealing more than a certain number of bits about the identity of the source. Our generalization leads to an IB-type tradeoff between the prediction of the target (generalized redundancy) and compression (leakage of information about the identity of the source). We refer to the resulting optimization problem as the redundancy bottleneck (RB) and to the manifold of optimal solutions at different points on the prediction/compression tradeoff as the RB curve. We also show that the RB prediction and compression terms can be decomposed into contributions from individual sources, giving rise to an individual RB curve for each source.
@@
Besides the intrinsic theoretical interest of unifying PID and the IB, our approach brings about several practical advantages. In particular, the RB curve offers a fine-grained analysis on PID redundancy, showing how redundant information emerges at various scales and across different sources. This fine-grained analysis can be used to uncover sets of redundant sources without performing intractable combinatorial optimization. Our approach also has numerical advantages. The original formulation of Blackwell redundancy was based on a difficult optimization problem that becomes infeasible for larger systems. By reformulating Blackwell redundancy as an IB-type problem, we are able to solve it efficiently using an iterative algorithm, even for larger systems (code available at https://github.com/artemyk/pid-as-ib, accessed 25-06-2024). Finally, the RB has some attractive formal properties. For instance, unlike the original Blackwell redundancy, the RB curve is continuous in the underlying probability distributions.
@@
This paper is organized as follows. In the next section, we provide the background on the IB, PID, and Blackwell redundancy. In Section III, we introduce the RB, illustrate it with several examples, and discuss its formal properties. In Section IV, we introduce an iterative algorithm to solve the RB optimization problem. We discuss the implications and possible future directions in Section V. All proofs are found in the Appendix.
@@ II Background
We begin by providing relevant background on the information bottleneck, partial information decomposition, and Blackwell redundancy.
@@ II.1 Information Bottleneck (IB)
The information bottleneck (IB) method provides a way to extract information that is present in one random variable $X$ that is relevant for predicting another target random variable $Y$ [6, 15, 16]. To do so, the IB posits a “bottleneck variable” $Q$ that obeys the Markov condition $Q-X-Y$. This Markov condition guarantees that $Q$ does not contain any information about $Y$ that is not found in $X$. The quality of any particular choice of bottleneck variable $Q$ is quantified via two mutual information terms: $I(X;Q)$, which decreases when $Q$ provides a more compressed representation of $X$, and $I(Y;Q)$, which increases when $Q$ allows a better prediction of $Y$. The IB method selects $Q$ to maximize prediction given a constraint on compression [16, 15, 17]:
$I_{\text{IB}}(R)=\max_{Q:Q-X-Y}I(Y;Q)\quad\textrm{where}\quad I(X;Q)\leq R.$ (1)
@@
In practice, the IB curve is usually explored by considering the Lagrangian relaxation of the constrained optimization problem (1):
$F_{\text{IB}}(\beta):=\max_{Q}I(Y;Q)-\frac{1}{\beta}I(X;Q)$ (2)
Here, $\beta\geq 0$ is a parameter that controls the tradeoff between compression cost (favored for $\beta\to 0$) and prediction benefit (favored for $\beta\to\infty$). The advantage of the Lagrangian formulation is that it avoids the non-linear constraint in Eq. (1). If the IB curve is strictly concave, then the two formulations (1) and (2) are equivalent, meaning that there is a one-to-one map between the solutions of both problems [18]. When the IB curve is not strictly concave, a modified objective such as the “squared Lagrangian” or “exponential Lagrangian” should be used instead; see Refs. [18, 19, 20] for more details.
Since the original proposal, many reformulations, generalizations, and variants of the IB have been developed [7]. Notable examples include the “conditional entropy bottleneck” (CEB) [13, 21], the “multi-view IB” [22], the “distributed IB” [23], as well as a large family of objectives called the “multivariate IB” [24]. All of these approaches consider some tradeoff between two information-theoretic terms: one that quantifies the prediction of target information that should be maximized and one that quantifies the compression of unwanted information that should be minimized. We refer to an optimization that involves a tradeoff between information-theoretic prediction and compression terms as an IB-type problem.
@@ II.2 Partial Information Decomposition
The PID considers how information about a target random variable $Y$ is distributed across a set of source random variables $X_{1},\dots,X_{n}$. One of the main goals of the PID is to quantify redundancy, the amount of shared information that is found in each of the individual sources. The notion of redundancy in PID was inspired by an analogy between sets and information that has re-appeared in various forms throughout the history of information theory [25, 26, 27, 28, 29, 30, 31]. Specifically, if the amount of information provided by each source is conceptualized as the size of a set, then the redundancy is conceptualized as the size of the intersection of those sets [1, 5, 4]. Until recently, however, this analogy was treated mostly as an informal source of intuition, rather than a formal methodology.
In a recent paper [4], we demonstrated that the terms of PID can be defined by formalizing this analogy to set theory. Recall that, in set theory, the intersection of sets $A_{1},\dots,A_{n}$ is defined as the largest set $B$ that is contained in each set $A_{s}$ for $s\in\{1...n\}$. Thus, the size of the intersection of finite sets $A_{1},\dots,A_{n}$ is
$\left|\bigcap_{s=1}^{n}A_{s}\right|=\max_{B}|B|\quad\textrm{where}\quad B\subseteq A_{s}\quad\forall s\in\{1\dots n\}.$
We showed that PID redundancy can be defined in a similar way: the redundancy between sources $X_{1},\dots,X_{n}$ is the maximum mutual information in any random variable $Q$ that is less informative about the target $Y$ than each individual source [4]:
$I_{\cap}^{\sqsubseteq}:=\max_{Q}I(Q;Y)\quad\textrm{where}\quad Q\sqsubseteq X_{s}\quad\forall s\in\{1\dots n\}.$ (3)
@@
We used a similar approach to define “union information”, which in turn leads to a principled measure of synergy [4]. Note that union information and redundancy are related algebraically but not numerically; in particular, unlike in set theory, the principle of inclusion–exclusion does not always hold.
As mentioned above, here, we focus entirely on redundancy and leave the exploration of connections between IB and union information/synergy for future work.
@@ II.3 Blackwell Redundancy
The Blackwell order is a preorder relation over “channels”, that is conditional distributions with full support. A channel $\kappa_{B|Y}$ is said to be less informative than $\kappa_{C|Y}$ in the sense of the Blackwell order if there exists some other channel $\kappa_{B|C}$ such that
@@
Throughout, we use the notation $\circ$ to indicate the composition of channels, as defined via matrix multiplication. For instance, $\kappa_{B|Y}=\kappa_{B|C}\circ\kappa_{C|Y}$ is equivalent to the statement $\kappa_{B|Y}(b|y)=\sum_{c}\kappa_{B|C}(b|c)\kappa_{C|Y}(c|y)$ for all $b$ and $y$. Eq. (4) implies that $\kappa_{B|Y}$ is less informative than $\kappa_{C|Y}$ if $\kappa_{B|Y}$ can be produced by downstream stochastic processing of the output of channel $\kappa_{C|Y}$. We use the notation
@@
to indicate that $\kappa_{B|Y}$ is less Blackwell-informative than $\kappa_{C|Y}$. The Blackwell order can also be defined over random variables rather than channels. Given a target random variable $Y$ with full support, random variable $B$ is said to be less Blackwell-informative than random variable $C$, written as
@@
when their corresponding conditional distributions obey the Blackwell relation, $p_{B|Y}\preceq p_{C|Y}$ [36]. It is not hard to verify that any random variable $B$ that is independent of $Y$ is lowest under the Blackwell order, obeying $B\preceq_{Y}C$ for all $C$.
@@
The Blackwell order plays a fundamental role in statistics, and it has an important operational characterization in decision theory [37, 36, 38]. Specifically, $p_{B|Y}\preceq p_{C|Y}$ if and only if access to channel $p_{C|Y}$ is better for every decision problem than access to channel $p_{B|Y}$. See Refs. [39, 4] for details of this operational characterization and Refs. [36, 39, 40, 4, 41, 42] for more discussion of the relation between the Blackwell order and the PID.
@@
Combining the Blackwell order (6) with Eq. (3) gives rise to Blackwell redundancy [4]. Blackwell redundancy, indicated here as $I_{\cap}$, is the maximal mutual information in any random variable that is less Blackwell-informative than each of the sources:
$I_{\cap}:=\max_{Q}\,I(Q;Y)\quad\textrm{where}\quad Q\preceq_{Y}X_{s}\;\forall s.$ (7)
The optimization is always well defined because the feasible set is not empty, given that any random variable $Q$ that is independent of $Y$ satisfies the constraints. (Note also that, for continuous-valued or countably infinite sources, $\max$ may need to be replaced by a $\sup$; see also Appendix Appendix.)
@@
$I_{\cap}$ has many attractive features as a measure of PID redundancy, and it overcomes several problems with previous approaches [4]. For instance, it can be defined for any number of sources, it uniquely satisfies a natural set of PID axioms, and it has fundamental statistical and operational interpretations. Statistically, it is the maximum information transmitted across any channel that can be produced by downstream processing of any one of the sources. Operationally, it is the maximum information that any random variable can have about $Y$ without being able to perform better on any decision problem than any one of the sources.
@@
As we showed [4], the optimization problem (7) can be formulated as the maximization of a convex function subject to a set of linear constraints. For a finite-dimensional system, the feasible set is a finite-dimensional polytope, and the maximum will lie on one of its extreme points; therefore, the optimization can be solved exactly by enumerating the vertices of the feasible set and choosing the best one [4]. However, this approach is limited to small systems, because the number of vertices of the feasible set can grow exponentially.
Finally, it may be argued that Blackwell redundancy is actually a measure of redundancy in the channels $p_{X_{1}|Y},\dots,p_{X_{n}|Y}$, rather than in the random variables $X_{1},\dots,X_{n}$. This is because the joint distribution over $(Y,X_{1},\dots,X_{n})$ is never explicitly invoked in the definition of $I_{\cap}$; in fact, any joint distribution is permitted as long as it is compatible with the correct marginals. (The same property holds for several other redundancy measures [Table 1, 4], and Ref. [39] even suggested this property as a requirement for any valid measure of PID redundancy.) In some cases, the joint distribution may not even exist, for instance when different sources represent mutually exclusive conditions. To use a neuroscience example, imagine that $p_{X_{1}|Y}$ and $p_{X_{2}|Y}$ represent the activity of some brain region $X$ in response to stimulus $Y$, measured either in younger ($p_{X_{1}|Y}$) or older ($p_{X_{2}|Y}$) subjects. Even though there is no joint distribution over $(Y,X_{1},X_{2})$ in this case, redundancy is still meaningful as the information about the stimulus that can be extracted from the brain activity of either age group. In the rest of this paper, we generally work within the channel-based interpretation of Blackwell redundancy.